McGilla Golf has decided to sell a new line of golf clubs. The clubs will sell for $830 per set and have a variable cost of $310 per set. The company has spent $215,000 for a marketing study that determined the company will sell 40,000 sets per year for seven years. The marketing study also determined that the company will lose sales of 9,000 sets of its high-priced clubs. The high-priced clubs sell at $1,260 and have variable costs of $730. The company will also increase sales of its cheap clubs by 9,000 sets. The cheap clubs sell for $490 and have variable costs of $120 per set. The fixed costs each year will be $10,750,000. The company has also spent $1,505,000 on research and development for the new clubs. The plant and equipment required will cost $25,000,000 and will be depreciated on a straight-line basis. The new clubs will also require an increase in net working capital of $1,445,000 that will be returned at the end of the project. The tax rate is 33 percent, and the cost of capital is 16 percent.

Suppose you feel that the values are accurate to within only ±9 percent. The best-case NPV is $ and worst-case NPV is $. (Do not include the dollar signs ($). Negative amount should be indicated by a minus sign. Round your answers to 2 decimal places. (e.g., 32.16)) (Hint: The price and variable costs for the two existing sets of clubs are known with certainty; only the sales gained or lost are uncertain.)
Answer #1
Cost of marketing study and research are sunk cost and not relevant for this analysis
BEST CASE SCENARIO:
A Sales Price per set                   904.70 (830*1.09)
B Variable Costs per set                   282.10 (310*0.91)
C=A-B Contribution margin per set                   622.60
D Quantity of sales per year                   43,600 (40000*1.09)
E=C*D Annual Contribution            27,145,360
F Contribution margin of high priced clubs                   530.00 (1260-730)
G Quantity of lost sales of high priced                     8,190 (9000*0.91)
H=F*G Loss of contribution per annum              4,340,700
I Contribution margin of Cheap clubs                   370.00 (490-120)
J Quantity of Increase in sales of Cheap clubs                     9,810 (9000*1.09)
K=J*I Addition to contribution per annum              3,629,700
L=E-H+K Total AnnualContributon            26,434,360
M Annual Fixed Costs              9,782,500 (10750000*0.91)
N=L-M Annual Profit (excluding depreciation)            16,651,860
P Cost of new equipment            22,750,000 (25000000*0.91)
Q=P/7 Depreciation per year              3,250,000
R=N-Q Before tax operating profit            13,401,860
S=R*(1-0.33) After tax operating profit(33% tax rate)              8,979,246
T=S+Q After tax annual operating cash flow            12,229,246
PV1 Present Value of annual operating cash flow            49,388,611 (Using PV function of excel with Rate=16%,Nper=7, Pmt=-T)
U Increase in net working capital              1,314,950 (1445000*0.91)
PV2 Present Value of returned working capital                 465,268 (1314950/(1.16^7)
NPV=PV1+PV2-P-U NET PRESENT VALUE (NPV) BEST CASE            25,788,929

WORST CASE SCENARIO Sales Price per set Variable Costs per set Contribution margin per set Quantity of sales per year Annual Contribution Contribution margin of high priced clubs Quantity of lost sales of high priced Loss of contribution per annum 755.30 (830*0.91) 337.90 (310 1.09) 417.40 36,400(40000 0.91) C-A-B 15,193,360 9,810 (9000 1.09) H-F*G Contribution margin of Cheap clubs Quantity of Increase in sales of Cheap clubs Addition to contribution per annum Total AnnualContributon Annual Fixed Costs Annual Profit (excluding depreciation) Cost of new equipment Depreciation per year Before tax operating profit After tax operating profit(33% tax rate) After tax annual operating cash flow Present Value of annual operating cash flow Increase in net working capital Present Value of returned working capital NET PRESENT VALUE (NP) BEST CASE 370.00(490-120) 8.190 (9000 0.91) KEJT L-E-H+K 13,024,360 11717.500 1(6107500001.09) N=L-M 27,250,000 3,892,857 (2.585.997 (1,732,618 (25000000*1.09) R-N-Q S R(1-0.33) T-S+0 PV1 (1,732,618 8,724,267 (Using PV function of excel with Rate: 16%,Nper-7, Pmt-T) 1575,050 1445000 1.09) PV2 NPV-PV1+PV2-P-U 557.299 1575050/1.16a7) 19.543.484